solve the following quadratic equations. Leave surds in your answers.

a) (x-3)^2-3=0

b) (3x-7)^2 =8


2a) x^4-5x^2+4=0

b) x-8=2squareroot x


1
Expert's answer
2021-10-04T12:17:58-0400

a) "(x-3)^{2}-3=0"

Add "3" to both sides:

"(x-3)^{2}-3+\textcolor{primary}{3}=0+\textcolor{primary}{3}"

Add like terms:

"(x-3)^{2}=3"

Take the square root for both sides:

"x-3=\pm \sqrt{3}"

Add "3" to both sides:

"x=3\pm \sqrt{3}"

The solutions are "x=3- \sqrt{3}" and "x=3+\sqrt{3}"

----------------------------------------------------------------------------------------------------------------------------------------

b)"(3x-7)^{2}=8"

Take the square root for both sides:

"3x-7=\pm \sqrt{8}"

Add "7" to both sides:

"3x=7\\pm \\sqrt{8}\n\u200b"

Divide both sides by "3":

"x=\frac{7\pm\sqrt{8}}{3}"

Simplify the expression "\sqrt{8}" as "2\sqrt{2}" :

"x=\frac{7\pm2\sqrt{2}}{3}"

----------------------------------------------------------------------------------------------------------------------------------------

2a)"x^4-5x^2+4=0"

Find two numbers whose product is "4" "the third term" and have the sum of "-5" "coefficient of the second term":

The numbers are "-4" and "-1", such that "-4 \times -1 =4" and "-4+(-1)=-5"

Now factor the expression:

"(x^2-1)(x^2-4)=0"

Using Zero-Product Property, it follows that:

"(x^2-1)=0" or "(x^2-4)=0"

Take the first equation:

"x^2-1=0"

Add "1" to both sides:

"x^2=1"

Take the square root for both sides:

"x=\pm1"

Take the second equation:

"x^2-4=0"

Add "4" to both sides:

"x^2=4"

Take the square root for both sides:

"x=\pm2"

Therefore, the solutions are "-1,1,-2, \text{and}~ 2"

----------------------------------------------------------------------------------------------------------------------------------------

b)"x-8=2\sqrt{x}"

Square both sides:

"(x-8)^2=(2\sqrt{x})^2"

Expand the square of the binomial on the left-hand side:

"x^2-16x+64=(2\sqrt{x})^2"

Square the right-hand side:

"x^2-16x+64=4x"

Subtract "4x" from both sides:

"x^2-16x+64-4x=4x-4x"

Add like terms:

"x^2-20x+64=0"

Write "-20x" as a difference:

"x^2-4x-16x+64=0"

Factor out "x" from the first two expressions:

"x(x-4)-16x+64=0"

Factor out "-16" from the second two expressions:

"x(x-4)-16(" "x-4)" "=0"

Factor out "(x-4)" from the expression:

"(x-4)(x-16)=0"

Using Zero-Product Property, it follows that:

"(x-4)=0~~\text{or}~~(x-16)=0"

"x=4~~\text{or}~~x=16"